# Planning for Robust Open-loop Pushing: Exploiting Quasi-static Belief Dynamics and Contact-informed Optimization

### v1

@@ Abstract
Non-prehensile manipulation such as pushing is typically subject to uncertain, non-smooth dynamics. However, modeling the uncertainty of the dynamics typically results in intractable belief dynamics, making data-efficient planning under uncertainty difficult. This article focuses on the problem of efficiently generating robust open-loop pushing plans. First, we investigate how the belief over object configurations propagates through quasi-static contact dynamics. We exploit the simplified dynamics to predict the variance of the object configuration without sampling from a perturbation distribution. In a sampling-based trajectory optimization algorithm, the gain of the variance is constrained in order to enforce robustness of the plan. Second, we propose an informed trajectory sampling mechanism for drawing robot trajectories that are likely to make contact with the object. This sampling mechanism is shown to significantly improve chances of finding robust solutions, especially when making-and-breaking contacts is required. We demonstrate that the proposed approach is able to synthesize bi-manual pushing trajectories, resulting in successful long-horizon pushing maneuvers without exteroceptive feedback such as vision or tactile feedback.
@@ 1 Introduction
Second, the mechanics of non-prehensile contacts are subject to uncertainty. This is due to the fact that dynamic effects such as friction between surfaces are difficult to predict, especially when manipulating objects without knowing their physical properties (Lynch and Mason (1995); Dogar and Srinivasa (2011); Ha et al. (2020)). Thus, when generating open-loop plans by assuming an accurate model of the dynamics, those plans are likely to fail as they do not take into account the underlying uncertainty. In Rodriguez (2021), this effect is described with an experiment of repeatedly picking-and-placing a queen chess piece, which fails if the queen is picked from the top. In contrast, picking the queen from the side stabilizes the repeated pick-and-place process. This raises the question of how to let robots deal with this uncertainty autonomously. Deploying fast feedback loops is an implicit approach to compensating for uncertain dynamics, such as done in model-predictive control schemes. Yet, modeling uncertainties and reasoning over an anticipated distribution of outcomes, i.e. a belief, is a more explicit way of coping with uncertain dynamics.
@@
In a real-world experiment we demonstrate that the proposed approach is able to synthesize robust bi-manual pushing trajectories in only a few seconds of planning time, consisting of long-horizon (up to 100 seconds) open-loop pushing maneuvers that include making and breaking contacts (cf. Fig. 1). The experimental results show that the interplay of i) and ii) is crucial for synthesizing robust behavior. To the best of our knowledge, the proposed algorithm is the first model-based planning approach that is able to synthesize robust plans for contact-rich manipulation without pre-defined manipulation primitives.
@@ 11       $\bm{u}\leftarrow$ concatenate$\left(\bm{u},\bm{u}^{*}_{0:H-1}\right)$
Figure 10: Setup for the real-world experiment. The two manipulators are considered as one bi-manual robot with two end-effectors, each equipped with a ball-shaped end-effector. The object (bottle, glass or can) is placed in front of the robot and the goal is to push the object along the target path (red circle). All objects were chosen due to their circular shape for the sake of a simple implementation of the quasi-static contact dynamics. The initial belief over the object position is Gaussian distributed with a mean equal to the initial object position and a covariance that reflects the uncertainty in the object detection. The contact dynamics are modeled probabilistically to reflect uncertainty. For the experiments with the can, we add additional perturbations by placing wooden cubes (red cubes) on the target path and by changing the center of mass to be off-center by placing a heavy tool in the can (yellow content of the can).
@@
7 Experiment: Bi-manual Pushing
7.1 Experimental Setup
We chose to validate the proposed algorithm with a bi-manual pushing task, consisting of two Franka robot arms that are equipped with ball-shaped end-effectors. Note that we treat the two robot arms as one bi-manual robot. We plan 2D trajectories for the ball-shaped end-effectors in a plane parallel to the table. The dynamics are modeled in this 2D plane, where the two robots are abstracted as two independent circles. The objects are abstracted as circles as well. Fig. 10 shows the experimental setup. Initially, the object (bottle, glass or can) is placed in front the bi-manual robot and the goal for the robot is to push the object along a circular target path. The belief over object positions is initialized with a Gaussian distribution. We furthermore modeled the noise in the contact dynamics with a Gaussian distribution and we tuned the covariance to capture the stochasticity in the contact dynamics. To prevent collisions between the two end-effectors, we include a collision constraint enforcing that the two end-effectors do not touch. We also add a constraint that prevents the robot from crossing its arms.
In order to evaluate qualitatively the effect of modeling the uncertainty in the contact dynamics, we compare the performance of the proposed algorithm with a deterministic baseline algorithm that is based on the proposed receding-horizon algorithm. For this, the planner assumes that the initial object position is known and it uses the nominal object dynamics from Sec. 4 as deterministic dynamics model. Consequently, we run the baseline without the cost and constraints on the variance gain.
@@ 7.2 Implementation
We implement the contact dynamics as in (6) for the special case of circular shapes in two dimensions. In order to approximate the initial belief via particles, we found that $N_{p}=20$ particles are sufficient to generate robust plans. For computing the CMA-ES updates to the Gaussian distribution over candidate trajectories, we use the Python package provided by the authors of Hansen (2016). We provide additional details on the cost design in the appendix. The planning algorithm was executed on a laptop with an Intel Core i9-14900HX CPU and 32GB of RAM. On average, generating a robust pushing plan for the full target path takes around seven seconds wall-clock time.
7.3 Qualitative Planning Results
We illustrate plans generated with the deterministic baseline and with the proposed receding-horizon BS-VP-STO algorithm in Fig. 11. Each plan is presented with 1000 stochastic rollouts of the object dynamics used for optimization. For an evaluation of the real-world performance of the plans, please refer to Sec. 7.5.
7.3.1 Deterministic Baseline.
In all plans generated with the deterministic baseline, the robot uses only one end-effector at a time for pushing the object. This is not surprising as the pushing progress, i.e. the mean control problem, is equally optimized when using one or two end-effectors. Thus, the baseline is not forced to discover the coordination between the two end-effectors and converges to the simpler solution, i.e. using one end-effector. When performing open-loop stochastic rollouts of the baseline plans, the object deviates from the planned trajectory after some time and thus the robot is not able to successfully push the object along the whole target path.
7.3.2 Receding-horizon BS-VP-STO.
In contrast, we observe that the robot uses both end-effectors to push the object along the target path when planning with the receding-horizon BS-VP-STO algorithm. Note that the strategy of using two end-effectors for pushing deliberately emerges from planning in belief space. The proposed algorithm discovers the use of two end-effectors by penalizing sampled robot trajectories that result in an increasing uncertainty about the object position, i.e. using one end-effector. At the beginning of the pushing maneuver, the robot performs an action that reduces uncertainty by placing its end-effectors such that they enclose the initial belief. This effectively brings the particles closer together, resulting in a decreasing variance. The robot then starts to make contact with its two end-effectors side-by-side to push the object along the target path with high probability. After pushing the object along the first half of the circular target path, the no-collision constraint between the two end-effectors forces the robot to break the contact and to find a new contact configuration with which it can continue pushing. Consequently, the robot has to move its left end-effector around the object without touching it. Afterwards, the robot makes contact again with the object and continues pushing until it reaches the initial position again. When performing open-loop stochastic rollouts of the robust plans, the robot has a high probability of being successful at pushing the object along the target path despite the perturbations and the lack of feedback.
7.4 Quantitative Planning Results & Ablation Studies
@@
7.4.1 Impact of the Number of BS-VP-STO Iterations.
Fig. 12 illustrates the statistics of success over the number of BS-VP-STO iterations. We observe that the success rate increases with the number of iterations $M$ and reaches around $100\%$ success rate with $M=4$ BS-VP-STO iterations. Note that running BS-VP-STO for one iteration poses a special case of the algorithm since no CMA-ES update is performed, resulting in only sampling an initial candidate population and picking the best performing candidate. This procedure in fact corresponds to the predictive sampling algorithm introduced in Howell et al. (2022).
@@
To evaluate the impact of the contact prior, we set the contact precision matrix in BS-VP-STO to $\bm{Q}_{\bm{q}}=\bm{0}$ (cf. (39)). This corresponds to uninformed trajectory samples as depicted in the left sub-figure of Fig. 6. In Fig. 12 we observe that the success rate of the algorithm without the contact prior is significantly lower than the success rate of the algorithm with the contact prior. This indicates that the contact prior is improving the efficiency of the planning algorithm to find robust manipulation actions in few iterations. Computing the contact prior requires a matrix inversion to compute $\bm{Q}_{\bm{q}}$ with dimensionality equivalent to the number of degrees of freedom. Since this operation has to be performed only once for a single-horizon, the computational overhead of the contact prior is negligible.
7.4.3 Scalability to Many Degrees of Freedom.
@@
Last, we evaluate the scalability of the proposed planning algorithm to many degrees of freedom by moving from planning in the 2D plane to planning in the joint space of the bi-manual robot, which has 14 degrees of freedom (seven degrees of freedom per robot arm). Coordinating all degrees of freedom adds complexity to the planning problem, while at the same time increasing the dimensionality of the search space. In Fig. 12 we observe that the success rate for a given number of iterations $M$ drops when increasing the complexity of the problem, requiring more iterations to discover the required coordination between the 14 joints. While the contact prior in joint space (cf. Sec. 5.2.3) imposes coordination between the seven joints of the individual arms, BS-VP-STO is still required to find joint trajectories such that the two end-effectors touch the object at the right place at the right time. The planning time per iteration is not significantly increased compared to planning for four degrees of freedom, since only an additional forward kinematics computation for the belief rollout is required. However, note that the contact dynamics are still modeled using the ball-shaped abstraction of the end-effector. Modeling the whole kinematic chain of the robot arms as contact geometries adds additional computational complexity to the planner.
7.5 Real-world Pushing Results
@@
Fig. 13 shows snapshots of the robot behavior planned with the receding-horizon BS-VP-STO algorithm. In addition, a full video of the experiment can be found in the supplementary material. We executed the planned robot trajectories for each of the three objects (bottle, glass, can), where we conducted five experiments for each object by placing the object at different initial positions. Out of 15 executed plans, the robot successfully pushed the object along the target path in 14 experiments. The robot failed to push the glass in one experiment, where the object was too far away from the mean initial position such the robot did not enclose the object during the initial uncertainty-reducing action.
@@
We furthermore evaluated the deterministic baseline planner with the same three objects, while placing the objects only at the expected location. In all three experiments, the actual object position deviated from the planned object position after a few pushes, resulting in the robot loosing contact with the object and thus failing to push the object along the target path. We show an example of the motion generated by the baseline in the supplementary video.

### v2

# Robust Pushing: Exploiting Quasi-static Belief Dynamics and Contact-informed Optimization

@@ Abstract
Non-prehensile manipulation such as pushing is typically subject to uncertain, non-smooth dynamics. However, modeling the uncertainty of the dynamics typically results in intractable belief dynamics, making data-efficient planning under uncertainty difficult. This article focuses on the problem of efficiently generating robust open-loop pushing plans. First, we investigate how the belief over object configurations propagates through quasi-static contact dynamics. We exploit the simplified dynamics to predict the variance of the object configuration without sampling from a perturbation distribution. In a sampling-based trajectory optimization algorithm, the gain of the variance is constrained in order to enforce robustness of the plan. Second, we propose an informed trajectory sampling mechanism for drawing robot trajectories that are likely to make contact with the object. This sampling mechanism is shown to significantly improve chances of finding robust solutions, especially when making-and-breaking contacts is required. We demonstrate that the proposed approach is able to synthesize bi-manual pushing trajectories, resulting in successful long-horizon pushing maneuvers without exteroceptive feedback such as vision or tactile feedback. We furthermore deploy the proposed approach in a model-predictive control scheme, demonstrating additional robustness against unmodeled perturbations.
@@ 1 Introduction
Second, the mechanics of non-prehensile contacts are subject to uncertainty. This is due to the fact that dynamic effects such as friction between surfaces are difficult to predict, especially when manipulating objects without knowing their physical properties, such as friction coefficients or mass distribution (Lynch and Mason (1995); Dogar and Srinivasa (2011); Ha et al. (2020)). Thus, when generating open-loop plans by assuming an accurate model of the dynamics, those plans are likely to fail as they do not take into account the underlying uncertainty. In Rodriguez (2021), this effect is described with an experiment of repeatedly picking-and-placing a queen chess piece, which fails if the queen is picked from the top. In contrast, picking the queen from the side stabilizes the repeated pick-and-place process. Uncertainty about physical parameters of objects is particularly unavoidable when robots interact with objects for the first time. This raises the question of how to let robots deal with this uncertainty autonomously. Deploying deterministic models in fast feedback loops is an implicit approach to compensating for uncertain dynamics, such as done in model-predictive control schemes. Yet they require accurate sensory measurements and achieve robustness against perturbations or modelling errors only by correcting the observed control errors. Modeling uncertainties and reasoning over an anticipated distribution of outcomes, i.e. a belief, is a more explicit way of coping with uncertain dynamics. Planning in belief-space enables open-loop execution of plans while still exhibiting robustness against modeled perturbations, thus not requiring sensors. Moreover, control errors are anticipated at planning time and can thus be prevented before they happen, potentially making manipulation under uncertainty more efficient. Closed-loop control in belief-space can then be achieved by continuously updating the belief based on observations and subsequent replanning, combining the benefits of reactivity with respect to unmodeled perturbations and the effectiveness of preventing control errors at planning time.
@@
In real-world experiments we demonstrate that the proposed approach is able to synthesize robust bi-manual pushing trajectories in only a few seconds of planning time, consisting of long-horizon (up to 100 seconds) open-loop pushing maneuvers that include making and breaking contacts (cf. Fig. 1). The experimental results show that the interplay of i) and ii) is crucial for synthesizing robust behavior. We furthermore show that the proposed planning algorithm is fast enough to be used in a model-predictive control loop, enabling a combination of reactivity and anticipatory robustness. To the best of our knowledge, the proposed algorithm is the first model-based planning approach that is able to synthesize robust plans for contact-rich manipulation without pre-defined manipulation primitives.
@@ 11       $\bm{u}\leftarrow$ concatenate$\left(\bm{u},\bm{u}^{*}_{0:H-1}\right)$
7 Experiments: Robust Pushing
This section presents robot experiments validating the theory and algorithmic approach developed in this article. We use objects that the robot has never interacted with before, with the geometry of the objects being the only information available. In all experiments, we compare the performance of the proposed approach against a baseline that only uses the nominal model of the contact dynamics without considering uncertainties. For this, the planner assumes that the initial object position is known and it uses the nominal object dynamics from Sec. 4 as a deterministic dynamics model. Consequently, we run the baseline without the cost and constraints on the variance gain.
@@ 7.1 Implementation
Figure 10: Open-loop single-hand pushing experiment: The task for the robot is to use the rectangular geometry of its hand (highlighted in green) to push the object with a circular geometry (highlighted in orange) into a target position without sensory feedback. The experiment consists of repeating the same pushing plan open-loop until the object diverged off the path such that the robot does not make contact anymore.
We implement the contact dynamics as in (6) for the special case of circular and rectangular shapes in two dimensions. In order to approximate the initial belief via particles, we found that $N_{p}=20$ particles are sufficient to generate robust plans. For computing the CMA-ES updates to the Gaussian distribution over candidate trajectories, we use the Python package provided by the authors of Hansen (2016). We provide additional details on the cost design in the appendix. The planning algorithm was executed on a laptop with an Intel Core i9-14900HX CPU and 32GB of RAM. On average, generating a robust pushing plan for the full target path takes around seven seconds wall-clock time.
7.2 Open-Loop Single-Hand Pushing
The first experiment takes Example 3 from simulation into the real world. As an end-effector, the robot uses the rectangular-shaped hand of the Franka robot to push the target object, without considering the fingers for contacts. Fig. 10 illustrates the experimental setup showing the initial robot configuration, the initial object position and the target object position. We compare trajectories of 2D positions and yaw angles of the hand generated by our approach (receding-horizon BS-VP-STO) and a baseline approach. The baseline optimizes for efficient trajectories assuming that the nominal contact dynamics accurately predict the object trajectory. To evaluate the robustness of generated plans, we execute the plan to let the robot push the object into the target position and use the same plan for pushing the object back to the initial position. Subsequently, the same plan is executed repeatedly open-loop with the object located where the previous execution ended. We expect that uncertainties in the contact dynamics will lead to deviations from the nominal model, resulting in the accumulation of control errors. Hence, we measure robustness by counting the number of successive runs until the robot loses the object. For a video showing qualitative results, see Extension 1. In the following, we report quantitative results.
7.2.1 Deterministic Baseline
As the baseline does not account for uncertainties in the contact dynamics, the resulting optimal trajectory consists of pushing in a straight line towards the target position while keeping the long side of the hand perpendicular to the pushing direction. We execute 10 experiments with the optimal baseline plan, resulting in an average of $6.8$ (min. $5$, max. $9$) successive runs until the robot loses the object. This is the result of deviations from the nominal model due to real-world perturbations such as imperfect friction surfaces and mass distributions, leading to an accumulation of errors in the object positions when pushing a circular object with a flat surface.
7.2.2 Receding-horizon BS-VP-STO
The proposed algorithm uses the same nominal model that was used in the baseline, while modeling additional uncertainty. The resulting optimal trajectory is executed in $10$ open-loop experiments without additional perturbations to compare the result with the baseline. All experiments have been stopped after $40$ successive runs as the system did not show any sign of accumulating errors. This indicates that the optimized trajectory actively controls the uncertainty in the object position by keeping track of the open-loop propagated belief.
7.3 Open-Loop Bi-manual Pushing
Figure 11: Open-loop bi-manual pushing experiment: The two manipulators are considered as one bi-manual robot with two end-effectors, each equipped with a ball-shaped end-effector. The object (bottle, glass or can) is placed in front of the robot and the goal is to push the object along the target path (red circle). All objects were chosen due to their circular shape for the sake of a simple implementation of the quasi-static contact dynamics. The initial belief over the object position is Gaussian distributed with a mean equal to the initial object position and a covariance that reflects the uncertainty in the object detection. The contact dynamics are modeled probabilistically to reflect uncertainty. For the experiments with the can, we add additional perturbations by placing wooden cubes (red cubes) on the target path and by changing the center of mass to be off-center by placing a heavy tool in the can (yellow content of the can).
We choose to validate the proposed algorithm with a bi-manual pushing task, consisting of two Franka robot arms that are equipped with ball-shaped end-effectors. Note that we treat the two robot arms as one bi-manual robot. We plan 2D trajectories for the ball-shaped end-effectors in a plane parallel to the table. The dynamics are modeled in this 2D plane, where the two robots are abstracted as two independent circles. The objects are abstracted as circles as well. Fig. 11 shows the experimental setup. Initially, the object (bottle, glass or can) is placed in front the bi-manual robot and the goal for the robot is to push the object along a circular target path. The belief over object positions is initialized with a Gaussian distribution. We furthermore modeled the noise in the contact dynamics with a Gaussian distribution and we tuned the covariance to capture the stochasticity in the contact dynamics. To prevent collisions between the two end-effectors, we include a collision constraint enforcing that the two end-effectors do not touch. We also add a constraint that prevents the robot from crossing its arms. Note that the time, location and number of contacts is subject to planning, i.e. we do not impose any heuristic that forces the robot to use both end-effectors for pushing. Instead, the proposed cost and constraint on the variance gain drives the optimization algorithm to find stabilizing contact configurations and sequences, such as using both end-effectors for pushing.
7.3.1 Qualitative Planning Results.
@@ 7.3 Open-Loop Bi-manual Pushing
We illustrate plans generated with the deterministic baseline and with the proposed receding-horizon BS-VP-STO algorithm in Fig. 12. Each plan is presented with 1000 stochastic rollouts of the object dynamics used for optimization. For an evaluation of the real-world performance of the plans, please refer to Sec. 7.3.3.
Deterministic Baseline. In all plans generated with the deterministic baseline, the robot uses only one end-effector at a time for pushing the object. This is not surprising as the pushing progress, i.e. the mean control problem, is equally optimized when using one or two end-effectors. Thus, the baseline is not forced to discover the coordination between the two end-effectors and converges to the simpler solution, i.e. using one end-effector. When performing open-loop stochastic rollouts of the baseline plans, the object deviates from the planned trajectory after some time and thus the robot is not able to successfully push the object along the whole target path.
Receding-horizon BS-VP-STO. In contrast, we observe that the robot uses both end-effectors to push the object along the target path when planning with the receding-horizon BS-VP-STO algorithm. Note that the strategy of using two end-effectors for pushing deliberately emerges from planning in belief space. The proposed algorithm discovers the use of two end-effectors by penalizing sampled robot trajectories that result in an increasing uncertainty about the object position, i.e. using one end-effector. At the beginning of the pushing maneuver, the robot performs an action that reduces uncertainty by placing its end-effectors such that they enclose the initial belief. This effectively brings the particles closer together, resulting in a decreasing variance. The robot then starts to make contact with its two end-effectors side-by-side to push the object along the target path with high probability. After pushing the object along the first half of the circular target path, the no-collision constraint between the two end-effectors forces the robot to break the contact and to find a new contact configuration with which it can continue pushing. Consequently, the robot has to move its left end-effector around the object without touching it. Afterwards, the robot makes contact again with the object and continues pushing until it reaches the initial position again. When performing open-loop stochastic rollouts of the robust plans, the robot has a high probability of being successful at pushing the object along the target path despite the perturbations and the lack of feedback.
7.3.2 Quantitative Planning Results & Ablation Studies.
@@
Impact of the Number of BS-VP-STO Iterations. Fig. 13 illustrates the statistics of success over the number of BS-VP-STO iterations. We observe that the success rate increases with the number of iterations $M$ and reaches around $100\%$ success rate with $M=4$ BS-VP-STO iterations. Note that running BS-VP-STO for one iteration poses a special case of the algorithm since no CMA-ES update is performed, resulting in only sampling an initial candidate population and picking the best performing candidate. This procedure in fact corresponds to the predictive sampling algorithm introduced in Howell et al. (2022).
@@
Impact of the Contact Prior. To evaluate the impact of the contact prior, we set the contact precision matrix in BS-VP-STO to $\bm{Q}_{\bm{q}}=\bm{0}$ (cf. (39)). This corresponds to uninformed trajectory samples as depicted in the left sub-figure of Fig. 6. In Fig. 13 we observe that the success rate of the algorithm without the contact prior is significantly lower than the success rate of the algorithm with the contact prior. This indicates that the contact prior is improving the efficiency of the planning algorithm to find robust manipulation actions in few iterations. Computing the contact prior requires a matrix inversion to compute $\bm{Q}_{\bm{q}}$ with dimensionality equivalent to the number of degrees of freedom. Since this operation has to be performed only once for a single-horizon, the computational overhead of the contact prior is negligible.
@@
Scalability to Many Degrees of Freedom. Last, we evaluate the scalability of the proposed planning algorithm to many degrees of freedom by moving from planning in the 2D plane to planning in the joint space of the bi-manual robot, which has 14 degrees of freedom (seven degrees of freedom per robot arm). Coordinating all degrees of freedom adds complexity to the planning problem, while at the same time increasing the dimensionality of the search space. In Fig. 13 we observe that the success rate for a given number of iterations $M$ drops when increasing the complexity of the problem, requiring more iterations to discover the required coordination between the 14 joints. While the contact prior in joint space (cf. Sec. 5.2.3) imposes coordination between the seven joints of the individual arms, BS-VP-STO is still required to find joint trajectories such that the two end-effectors touch the object at the right place at the right time. The planning time per iteration is not significantly increased compared to planning for four degrees of freedom, since only an additional forward kinematics computation for the belief rollout is required. However, note that the contact dynamics are still modeled using the ball-shaped abstraction of the end-effector. Modeling the whole kinematic chain of the robot arms as contact geometries adds additional computational complexity to the planner.
7.3.3 Real-world Pushing Results
@@
Fig. 14 shows snapshots of the robot behavior planned with the receding-horizon BS-VP-STO algorithm. In addition, a full video of the experiment can be found in Extension 2. We executed the planned robot trajectories for each of the three objects (bottle, glass, can), where we conducted five experiments for each object by placing the object at different initial positions. Out of 15 executed plans, the robot successfully pushed the object along the target path in 14 experiments. The robot failed to push the glass in one experiment, where the object was too far away from the mean initial position such the robot did not enclose the object during the initial uncertainty-reducing action.
@@
We furthermore evaluated the deterministic baseline planner with the same three objects, while placing the objects only at the expected location. In all three experiments, the actual object position deviated from the planned object position after a few pushes, resulting in the robot loosing contact with the object and thus failing to push the object along the target path. We show a video example of the motion generated by the baseline in Extension 2.
7.4 Closed-Loop Bi-manual Pushing
Figure 15: Closed-loop bi-manual pushing experiment: The left image shows the external perturbations applied by manually moving the object. A camera continuously provides noisy observations of the object position to the closed-loop controller. The right image shows how the proposed approach is able to generate robust plans in real-time, enabling uncertainty-aware model-predictive control loops.
Last, we show that receding-horizon BS-VP-STO can be used in a closed control loop, gaining additional robustness against out-of-distribution disturbances. We use the same bi-manual robot setup as in Sec. 7.3. However, instead of pushing an object along a target path, the task is to push the object to the center of the table while its position is perturbed by a human. In addition, noisy observations of the object position are provided by a camera for closing the loop. We deploy a particle filter for continuously updating the belief based on both the stochastic rollout and noisy observations of the object as described in Sec. 6. We qualitatively compare the resulting behavior of the robot when being controlled with receding-horizon BS-VP-STO against the behavior when using the deterministic nominal model in a model-predictive controller. Both controllers run at a control rate of $5$ Hz with $M=6$ BS-VP-STO iterations per control step. Fig. 15 illustrates the external perturbations applied to the object and the resulting robot behavior when controlled with our proposed robust approach. A video of the resulting behavior of both control approaches can be found in Extension 3.
We observe that the robot robustly pushes the object back to the center of the table using both end-effectors when controlled with the proposed approach. The additional state estimation enables the robot to also react to out-of-distribution perturbations, re-generating robust plans given noisy measurement updates. For the deterministic baseline, we observe that the continuous feedback enables the robot to maintain contact with the object and to generate consistent pushes. However, we observe that the robot only uses one end-effector for pushing as also observed for the open-loop experiment in Sec. 7.3. In contrast to our approach, this leads to larger control errors during pushing that need to be corrected, resulting in the deterministic baseline taking more time to accurately bring the object back to the center of the table.